xXMarcelieXx:

can anyone help me with change of variables for calc 3

Professional:

no bieja

xXMarcelieXx:

uy biejaa lmao

Professional:

looool

xXMarcelieXx:

@Aeon

xXMarcelieXx:

help me with ellipses pls

Vocaloid:

@sillybilly123 whenever you get a chance

xXMarcelieXx:

xXMarcelieXx:

i need help with that one too

sillybilly123:

for: \(\left( \dfrac{x}{a}\right)^2 + \left( \dfrac{y}{b}\right)^2 = 1\) let \(u = \dfrac{x}{a}, v = \dfrac{y}{b}\), that's the unit circle, plus be sure to add in the Jacobian or for line integrals, which ALWAYS have to be expressed in a single parameter, let \(x = a \cos\theta, t = b \sin \theta\), so \(\theta\) is THAT parameter but quite hard to help w/o a concrete example

xXMarcelieXx:

oh okay

sillybilly123:

have you an actual question?!

sillybilly123:

you won't get the real answer to this kinda question inside Calc 3, you need to look at the metric tensor i think.

xXMarcelieXx:

i got it ;p

xXMarcelieXx:

or idk if i did it right lmao one sec

sillybilly123:

super duper 🤗 if you have a specific qu that irks you, post it here.